# Direct Learning of Mesh and Appearance via 3D Gaussian Splatting

### v2

@@ Abstract
Accurately reconstructing a 3D scene including explicit geometry information is both attractive and challenging. Geometry reconstruction can benefit from incorporating differentiable appearance models, such as Neural Radiance Fields and 3D Gaussian Splatting (3DGS). However, existing methods encounter efficiency issues due to indirect geometry learning and the paradigm of separately modeling geometry and surface appearance. In this work, we propose a learnable scene model that incorporates 3DGS with an explicit geometry representation, namely a mesh. Our model learns the mesh and appearance in an end-to-end manner, where we bind 3D Gaussians to the mesh faces and perform differentiable rendering of 3DGS to obtain photometric supervision. The model creates an effective information pathway to supervise the learning of both 3DGS and mesh. Experimental results demonstrate that the learned scene model not only achieves state-of-the-art efficiency and rendering quality but also supports manipulation using the explicit mesh. In addition, our model has a unique advantage in adapting to scene updates, thanks to the end-to-end learning of both mesh and appearance.
Index Terms:
@@
I Introduction
Mesh is often the preferred choice in numerous geometry processing scenarios, as it carry explicit surface information. When building a 3D scene model that includes a mesh, it is challenging to directly learn from visual observation due to the complicated pipeline required to produce the synthetic views. Recently, neural surface reconstruction methods [28, 32] have been developed to represent scene geometry as an implicit signed distance function (SDF). These methods establish a connection with Neural Radiance Fields (NeRFs) [17], enabling effective image-based supervision for geometry learning. However, the geometry is learned indirectly, as it involves converting the SDF to a radiance field. Additionally, NeRF-based learning requires expensive ray-marching in the volumetric rendering. These factors lead to efficiency challenges in learning geometry and appearance.
3D Gaussian Splatting (3DGS) [14] has recently gained popularity due to its ability to render photorealistic images significantly faster than NeRFs. 3DGS employs an explicit representation called Gaussians, which represent anisotropic ellipsoids in 3D space. Gaussians enable efficient rendering due to their explicit nature and the highly parallelized splatting procedure. Despite the advantages in rendering quality and speed, it is not straightforward to reconstruct object surfaces in the traditional form of Gaussians. This is because the 3DGS model imposes no explicit constraints on the structure of the Gaussians. As illustrated in Fig. 1, Gaussians that result in visually appealing rendering may be geometrically inaccurate – they can be inconsistent with any 2D manifold. Recently, some methods have incorporated regularization into the learning process of 3DGS. NeuSG [3] refines Gaussians using surface normals estimated by a jointly trained NeuS model [28], which notably increases the training time. SuGaR [10] encourages Gaussians to align with the surface and subsequently converts them into a point cloud for Poisson Surface Reconstruction. Once a coarse mesh is extracted, an additional stage is required to learn surface appearance from scratch using a new set of Gaussians. This two-stage pipeline may introduce efficiency problems. It is worth noting that similar issues are encountered in existing downstream applications [20, 12, 8, 6], which perform meshing and Gaussian training separately. These challenges motivate us to design a more effective information pathway to supervise the learning of scene representations.
Figure 1: The Gaussians and the underlying surface did not align well in the original 3DGS [14], whereas our hybrid representation explicitly restricts the Gaussians to the mesh faces. Our method benefits from the high-quality rendering of 3DGS and the controllability of a mesh.
@@ I Introduction
Extensive experiments show that our method achieves state-of-the-art rendering accuracy and produces high-quality meshes in indoor and outdoor scenes. Furthermore, it reduces total reconstruction time by 50% compared to recent methods and is over 10x faster than methods that rely on Neural SDFs and NeRFs [15, 28]. We also demonstrate that our method has a unique advantage in adapting to scene updates compared to other two-stage methods [10, 27].
@@
II Related Work
II-A Scene Reconstruction via Radiance Field
Neural Radiance Field (NeRF) [17] is a significant advancement in 3D reconstruction. It uses neural networks to model the radiance field and yield high-fidelity images via volumetric rendering. Despite its outstanding rendering quality, NeRF requires a long time for training and rendering due to the large number of samples needed to query the neural network. Recent efforts to accelerate NeRF have employed explicit grids [7] or hybrids of grids and small MLPs [24, 18]. However, these computational accelerations might compromise image quality. In addition to improving speed, some works explore scene manipulation and composition within NeRF [31], but these are still not convenient for downstream applications. This has led to another series of works [26, 28] that connect NeRF with a mesh, which will be discussed in the next section.
Recently, 3D Gaussian Splatting (3DGS) [14] employs anisotropic Gaussians to model an explicit radiance field, achieving real-time rendering and enhanced quality. The explicit nature of 3DGS makes it suitable for various domains such as SLAM [13] and AIGC [5], where efficiency and controllability are crucial. However, directly manipulating 3DGS is challenging due to the vast number of Gaussian components involved. Hence, existing applications like Avatars [20], VR [12], and physics simulations [8, 6] leverage mesh guidance to manipulate the scene more effectively.
II-B Image-based Surface Reconstruction
Reconstructing surfaces from images is a fundamental task in computer vision. Traditional methods, such as Multi-view Stereo (MVS) [2, 22], estimate depth maps or extract voxel grids based on correspondences between images. However, their quality is heavily dependent on the accuracy of image matching, and the use of voxel grids is often constrained by their cubic memory requirements. Recent advancements [19, 23] utilize differentiable isosurface extraction to directly optimize surface meshes under image supervision. Nonetheless, these methods are mainly for single objects due to the substantial memory demands of larger scenes.
Methods represented by NeuS [28, 29, 15] introduce a hybrid representation by transforming a Signed Distance Function (SDF) into a density field, adopting the learning pipeline of NeRFs. These techniques can model view-dependent appearances and also allow for the extraction of meshes from the SDF, making them suitable for downstream applications. To enable real-time rendering, some approaches [32, 21, 26] bake the appearance from a trained NeRF into the mesh textures. However, they typically require extensive training times and might reduce rendering quality.
@@ II-B Image-based Surface Reconstruction
III Preliminary
III-A 3D Gaussian Splatting
@@ III Preliminary
III-B Signed Distance Function
@@ IV Method
IV-A Signed Distance Function and Differentiable Mesh Computation
@@
Visible Node Optimization
During each training iteration, only the scene elements within a single camera’s viewing frustum are learning. Therefore, we extract mesh faces only from the visible grid nodes in $\{\boldsymbol{x}_{n}\}_{n=1}^{N}$ using the camera’s intrinsic and extrinsic parameters.
Coarse-to-fine Grid
Starting with a high-resolution grid can lead to redundant faces, causing unstable learning and memory issues. To address this, we begin with a low-resolution grid and progressively increase the resolution during training. The resolution is enhanced by linearly interpolating between the nodes of the current grid and the refined grid.
@@
Figure 3: Sub-figure (a) shows using $K=1,3,6$ Gaussians to represent the appearance of a triangle face, (b) defines a local coordinate frame, and (c) illustrates applying linear transformation $\boldsymbol{M}$ to let Gaussians adapt to the irregular triangle. See Sec. IV-B for more details.
IV-B Gaussians Consistent with Scene Geometry
@@
IV-C Gaussian Appearance Model
@@ V-A1 Initialization
V-A2 Joint Learning of Mesh and Gaussians
@@ V-B Datasets and Metrics
TABLE I: Per-scene quantitative comparisons on NeRF-Synthetic dataset [17].
@@
TABLE II: Quantitative comparisons on Mip-NeRF360 dataset [1].
@@
V-C Novel View Synthesis
We first evaluate our method on the novel view synthesis task. Tabs. I and II present the quantitative results on NeRF-Synthetic and Mip-NeRF360 datasets. To ensure a fair comparison, we employed a white background for the NeRF-Synthetic dataset. Our approach is compared against various recent methods, including those with reconstructed mesh and those without.
@@ V-C Novel View Synthesis
TABLE III: Quantitative Evaluation of Mesh Quality and Efficiency on the NeRF-Synthetic [17] dataset.
@@ V-D Surface Reconstruction
For a quantitative evaluation of the meshes, we compare the reconstructed mesh to the ground truth from the NeRF-Synthetic dataset using the Chamfer Distance (CD) metric. The computation method follows NeRF2Mesh [26], where 2.5M points are sampled from the surfaces of both the ground truth and the reconstructed object through ray casting. For a more comprehensive assessment, Tab. III presents the CD results, the number of faces, the rendering PSNR, and the training time. The results demonstrate that our reconstructed meshes achieve both high quality and training efficiency. Additionally, our Gaussian-based appearance model enables high-quality rendering even with fewer faces compared to NeRF2Mesh.
@@
V-E Object Deformation by Hybrid Representation
As discussed in our introduction section, the hybrid representation not only benefits from the high-quality and efficient rendering of Gaussians but also offers the convenience of manipulation through the mesh. To demonstrate the potential applications of this work, we selected object deformation tasks for testing. Specifically, we deform the learned mesh using four common mesh modifiers (Twist, Stretch, Bend, and Taper) in Blender and correspondingly adjust the bound Gaussians using the technique described in Sec. IV-B. We then use the Gaussian-based rasterization to render images of the deformed object, as shown in Fig. 7. More advanced operations, as explored in [12, 9], are also compatible with the representation we have obtained.
V-F Method Analysis
@@
TABLE IV: Evaluation on Efficiency. We report the training time and PSNR metric across different methods.
@@
TABLE V: Ablation studies on adaptive covariance and refinement stage. We report the average PSNR across all scenes of the NeRF-Synthetic [17] and Mip-NeRF360 [1] datasets.
Adaptive Cov. Refine NeRF-Synthetic [17] Mip-NeRF360 [1]
@@
TABLE VI: Ablation Study on the Number of Gaussians per Face. This table shows the PSNR results on the Mip-NeRF360 [1] dataset.
PSNR Stage-2
#.GS/face=1 #.GS/face=3 #.GS/face=6
Stage-1 #.GS/face=1 26.80 27.12 27.22
#.GS/face=3 26.89 27.21 27.54
‘Stage-1’: Joint learning of mesh and Gaussian appearance.
‘Stage-2’: Refinement stage that fixes mesh faces.
Note: The table does not include #.GS/$\geq$6 for Stage-1 due to OOM issues in some scenes.
V-F1 Efficiency
We assess our method’s efficiency by comparing the average time taken to reconstruct scenes from the NeRF-Synthetic and Mip-NeRF360 datasets. Both our method and the baseline methods are tested under identical hardware conditions. The results in Tab. IV reveal that our method is not only faster but also provides superior rendering quality.
V-F2 Adaptability in Scene Updates
@@
V-F3 Ablation Studies
In this section, we examine the impact of three factors: 1) the adaptive Gaussian covariance, 2) the refinement stage, and 3) the number of Gaussians per face. To evaluate the effect of adaptive Gaussian covariance, we set $\boldsymbol{M}$ in Eq. (10) to be an identity matrix. Tab. V shows that both adaptive Gaussian covariance and the refinement stage significantly enhance rendering quality.
The choice of the number of Gaussians per face involves a trade-off between computational consumption and performance. As shown in Tab. VI, increasing the number of Gaussians per face benefits rendering quality. To control the memory footprint and training time within an acceptable range while still achieving excellent rendering quality, we choose 3 Gaussians per face during the joint learning of mesh and Gaussian appearance, but use 6 Gaussians for the refinement stage.
VI Conclusion

### v3

@@ Abstract
Accurately reconstructing a 3D scene including explicit geometry information is both attractive and challenging. Geometry reconstruction can benefit from incorporating differentiable appearance models, such as Neural Radiance Fields and 3D Gaussian Splatting (3DGS). However, existing methods encounter efficiency issues due to indirect geometry learning and the paradigm of separately modeling geometry and surface appearance. In this work, we propose a learnable scene model that incorporates 3DGS with an explicit geometry representation, namely a mesh. Our model learns the mesh and appearance in an end-to-end manner, where we bind 3D Gaussians to the mesh faces and perform differentiable rendering of 3DGS to obtain photometric supervision. The model creates an effective information pathway to supervise the learning of both 3DGS and mesh. Experimental results demonstrate that the learned scene model not only improves efficiency and rendering quality but also enables manipulation via the explicit mesh. In addition, our model has a unique advantage in adapting to scene updates, thanks to the end-to-end learning of both mesh and appearance. The code is available at https://github.com/Cenbylin/DMGS.
keywords:
@@
1 Introduction
Mesh is often the preferred choice in numerous geometry processing scenarios, as it carry explicit surface information. When building a 3D scene model that includes a mesh, it is challenging to directly learn from visual observation due to the complicated pipeline required to produce the synthetic views. Recently, neural surface reconstruction methods [1, 2, 3] have been developed to represent scene geometry as an implicit signed distance function (SDF). These methods establish a connection with Neural Radiance Fields (NeRFs) [4], enabling effective image-based supervision for geometry learning. However, the geometry is learned indirectly, as it involves converting the SDF to a radiance field. Additionally, NeRF-based learning requires expensive ray-marching in the volumetric rendering. These factors lead to efficiency challenges in learning geometry and appearance.
3D Gaussian Splatting (3DGS) [5] has recently gained popularity due to its ability to render photorealistic images significantly faster than NeRFs. 3DGS employs an explicit representation called Gaussians, which represent anisotropic ellipsoids in 3D space. Gaussians enable efficient rendering due to their explicit nature and the highly parallelized splatting procedure. Despite the advantages in rendering quality and speed, it is not straightforward to reconstruct object surfaces in the traditional form of Gaussians. This is because the 3DGS model imposes no explicit constraints on the structure of the Gaussians. As illustrated in Fig. 1, Gaussians that result in visually appealing rendering may be geometrically inaccurate: they can be inconsistent with any 2D manifold. Recently, some methods have incorporated regularization into the learning process of 3DGS. NeuSG [6] refines Gaussians using surface normals estimated by a jointly trained NeuS model [1], which notably increases the training time. SuGaR [7] encourages Gaussians to align with the surface and subsequently converts them into a point cloud for Poisson Surface Reconstruction. Once a coarse mesh is extracted, an additional stage is required to learn surface appearance from scratch using a new set of Gaussians. This two-stage pipeline may introduce efficiency problems. It is worth noting that similar issues are encountered in existing downstream applications [8, 9, 10, 11], which perform meshing and Gaussian training separately. These challenges motivate us to design a more effective information pathway to supervise the learning of scene representations.
Figure 1: The Gaussians and the underlying surface did not align well in the original 3DGS [5], whereas our hybrid representation explicitly restricts the Gaussians to the mesh faces. Our method benefits from the high-quality rendering of 3DGS and the explicit surface structure provided by a mesh.
@@ 1 Introduction
Extensive experiments show that our method achieves superior rendering accuracy and produces high-quality meshes in indoor and outdoor scenes. Moreover, it reduces total reconstruction time by 50% compared to recent methods and is over 10x faster than methods that rely on Neural SDFs and NeRFs [14, 1]. Additionally, we show that our method offers a unique advantage in adapting to scene updates, outperforming existing two-stage methods [7, 15].
@@
2 Related Work
2.1 Scene Reconstruction via Radiance Field
Neural Radiance Field (NeRF) [4] is a significant advancement in 3D reconstruction. It uses neural networks to model the radiance field and yield high-fidelity images via volumetric rendering. Despite its outstanding rendering quality, NeRF requires a long time for training and rendering due to the large number of samples needed to query the neural network. Recent efforts to accelerate NeRF have employed explicit grids [16] or hybrids of grids and small MLPs [17, 18, 19]. However, these computational accelerations might compromise image quality. In addition to improving speed, some works explore scene manipulation and composition within NeRF [20], but these are still not convenient for downstream applications. This has led to another series of works [21, 1] that connect NeRF with a mesh, which will be discussed in the next section.
Recently, 3D Gaussian Splatting (3DGS) [5] employs anisotropic Gaussians to model an explicit radiance field, achieving real-time rendering and enhanced quality. The explicit nature of 3DGS makes it suitable for various domains such as scene understanding [22, 23] and AIGC [24], where efficiency and controllability are crucial. However, directly manipulating 3DGS is challenging due to the vast number of Gaussian components involved. Hence, existing applications like Avatars [8], VR [9], and physics simulations [10, 11] leverage mesh guidance to manipulate the scene more effectively.
2.2 Image-based Surface Reconstruction
Reconstructing surfaces from images is a fundamental task in computer vision. Traditional methods, such as Multi-view Stereo (MVS) [25, 26, 27], estimate depth maps or extract voxel grids based on correspondences between images. However, their quality is heavily dependent on the accuracy of image matching, and the use of voxel grids is often constrained by their cubic memory requirements. Recent advancements [12, 13] utilize differentiable isosurface extraction to directly optimize surface meshes under image supervision. Nonetheless, these methods are mainly for single objects due to the substantial memory demands of larger scenes.
Another promising direction, as demonstrated by NeuS [1, 28, 14], adopts a hybrid representation by transforming a Signed Distance Function (SDF) into a density field, adopting the learning pipeline of NeRFs. These techniques can model view-dependent appearances and also allow for the extraction of meshes from the SDF, making them suitable for downstream applications. To enable real-time rendering, some approaches [2, 29, 21] bake the appearance from a trained NeRF into the mesh textures. However, they typically require extensive training times and might reduce rendering quality.
@@ 2.2 Image-based Surface Reconstruction
3 Preliminary
3.1 3D Gaussian Splatting
@@ 3 Preliminary
3.2 Signed Distance Function
@@ 4 Method
4.1 Signed Distance Function and Differentiable Mesh Computation
@@
[…]
Visible Node Optimization During each training iteration, only the scene elements within a single camera’s viewing frustum are learning. Therefore, we extract mesh faces only from the visible grid nodes in $\{\boldsymbol{x}_{n}\}_{n=1}^{N}$ using the camera’s intrinsic and extrinsic parameters.
[…]
Coarse-to-fine Grid Starting with a high-resolution grid can lead to redundant faces, causing unstable learning and memory issues. To address this, we begin with a low-resolution grid and progressively increase the resolution during training. The resolution is enhanced by linearly interpolating between the nodes of the current grid and the refined grid.
@@
Figure 3: Sub-figure (a) shows using $K=1,3,6$ Gaussians to represent the appearance of a triangle face, (b) defines a local coordinate frame, and (c) illustrates applying linear transformation $\boldsymbol{M}$ to let Gaussians adapt to the irregular triangle. See Sec. 4.2 for more details.
4.2 Gaussians Consistent with Scene Geometry
@@
4.3 Gaussian Appearance Model
@@ 5.1.1 Initialization
5.1.2 Joint Learning of Mesh and Gaussians
@@ 5.2 Datasets and Metrics
Table 1: Per-scene quantitative comparisons on NeRF-Synthetic dataset [4].
@@
Table 2: Quantitative comparisons on Mip-NeRF360 dataset [36].
@@
5.3 Novel View Synthesis
We first evaluate our method on the novel view synthesis task. Tabs. 1 and 2 present the quantitative results on NeRF-Synthetic and Mip-NeRF360 datasets. To ensure a fair comparison, we employed a white background for the NeRF-Synthetic dataset. Our approach is compared against various recent methods, including those with reconstructed mesh and those without.
@@ 5.3 Novel View Synthesis
Table 3: Quantitative Evaluation of Mesh Quality and Efficiency on the NeRF-Synthetic [4] dataset.
@@ 5.4 Surface Reconstruction
For a quantitative evaluation of the meshes, we compare the reconstructed mesh to the ground truth from the NeRF-Synthetic dataset using the Chamfer Distance (CD) metric. The computation method follows NeRF2Mesh [21], where 2.5M points are sampled from the surfaces of both the ground truth and the reconstructed object through ray casting. For a more comprehensive assessment, Tab. 3 presents the CD results, the number of faces, the rendering PSNR, and the training time. The results demonstrate that our reconstructed meshes achieve both high quality and training efficiency. Additionally, our Gaussian-based appearance model enables high-quality rendering even with fewer faces compared to NeRF2Mesh.
@@
5.5 Object Deformation by Hybrid Representation
As discussed in our introduction section, the hybrid representation not only benefits from the high-quality and efficient rendering of Gaussians but also offers the convenience of manipulation through the mesh. To demonstrate the potential applications of this work, we selected object deformation tasks for testing. Specifically, we deform the learned mesh using four common mesh modifiers (Twist, Stretch, Bend, and Taper) in Blender and correspondingly adjust the bound Gaussians using the technique described in Sec. 4.2. We then use the Gaussian-based rasterization to render images of the deformed object, as shown in Fig. 7. More advanced operations, as explored in [9, 37], are also compatible with the representation we have obtained.
5.6 Method Analysis
@@
Table 4: Evaluation on Efficiency. We report the training time and PSNR metric across different methods.
@@
Table 5: Ablation studies on adaptive covariance and refinement stage. We report the average PSNR across all scenes of the NeRF-Synthetic [4] and Mip-NeRF360 [36] datasets.
Adaptive Cov. Refine NeRF-Synthetic Mip-NeRF360
@@
5.6.1 Efficiency
We assess our method’s efficiency by comparing the average time taken to reconstruct scenes from the NeRF-Synthetic and Mip-NeRF360 datasets. Both our method and the baseline methods are tested under identical hardware conditions. The results in Tab. 4 reveal that our method is not only faster but also provides superior rendering quality.
5.6.2 Adaptability in Scene Updates
@@
5.6.3 Ablation Studies
In this section, we examine the impact of three factors: 1) the adaptive Gaussian covariance, 2) the refinement stage, and 3) the number of Gaussians per face. To evaluate the effect of adaptive Gaussian covariance, we set $\boldsymbol{M}$ in Eq. (10) to be an identity matrix. Tab. 5 shows that both adaptive Gaussian covariance and the refinement stage significantly enhance rendering quality.
The choice of the number of Gaussians per face involves a trade-off between computational consumption and performance. As shown in Tab. 6, increasing the number of Gaussians per face benefits rendering quality. To control the memory footprint and training time within an acceptable range while still achieving excellent rendering quality, we choose 3 Gaussians per face during the joint learning of mesh and Gaussian appearance, but use 6 Gaussians for the refinement stage.
Table 6: Ablation Study on the Number of Gaussians per Face. This table shows the PSNR results on the Mip-NeRF360 [36] dataset.
6 Conclusion
@@
[…]
Data availability
The datasets analyzed during the current study are available in these repositories: NeRF-Synthetic https://github.com/bmild/nerf; Mip-NeRF360 https://jonbarron.info/mipnerf360.
Appendix A
Table 7: Per-scene quantitative results on Mip-NeRF360 dataset [36].
Metric Bicycle Bonsai Counter Garden Kitchen Room Stump Mean
PSNR 24.18 31.47 27.16 26.41 28.96 29.71 24.91 27.54
SSIM 0.722 0.945 0.888 0.829 0.898 0.908 0.712 0.843
LPIPS 0.260 0.175 0.212 0.161 0.162 0.221 0.291 0.212
A.1 Details for SDF Grid Optimization
This section includes the methods and configurations used to mitigate the storage and computational demands of the SDF grid:
[…]
Visible Node Optimization For the given grid nodes $\{\boldsymbol{x_{n}}\}_{n=1}^{N}$, the visible subset is calculated by:
$\{\boldsymbol{\hat{x}_{m}}\}:=\left\{\boldsymbol{x_{n}}\big{|}\text{NDC}(\boldsymbol{x_{n}})\in[-1,1]^{3}\right\},$ (12)
where $\text{NDC}(\boldsymbol{x_{n}})$ is the transformation from world-space coordinates $\boldsymbol{x_{n}}$ to their normalized device coordinates (NDC) via the projection matrix. The marching algorithm is only applied to cells whose eight corner nodes are within $\{\boldsymbol{\hat{x}_{m}}\}$.
[…]
Coarse-to-fine Grid Given the bounding box (BBox) with size ($L_{x}$, $L_{y}$, $L_{z}$) that tightly encloses the foreground region $U$, and an expected cell count $C$, the cubic cell size of the SDF grid is set to $(L_{x}/s)\times(L_{y}/s)\times(L_{z}/s)$, where $s=\sqrt[3]{L_{x}\cdot L_{y}\cdot L_{z}/C}$. During optimization, we perform 4 coarse-to-fine operations, incrementally increasing the cell count by 1.5×1.5$\times$1.5 × each step until the target count $C$ is reached.
A.2 Configurations of Face Gaussians
As discussed in Sec. 4.2, each face is associated with $K$ Gaussians. The choice of $K$ influences the barycentric coordinates of the Gaussian centers and the radius $r$ employed in the covariance matrix $\boldsymbol{\Sigma}_{e}$ in Eq. (10). For $K=3$, we use
$\boldsymbol{\xi}_{1}$ $=\left[(3-\sqrt{3})/6,(3-\sqrt{3})/6,\sqrt{3}/3\right],$
$\boldsymbol{\xi}_{2}$ $=\left[(3-\sqrt{3})/6,\sqrt{3}/3,(3-\sqrt{3})/6\right],$
$\boldsymbol{\xi}_{3}$ $=\left[\sqrt{3}/3,(3-\sqrt{3})/6,(3-\sqrt{3})/6\right],$
and radius $r=\|\boldsymbol{v}_{1}-\boldsymbol{v}_{2}\|/(2\sqrt{3}+2)$. For $K=6$, we use
$\boldsymbol{\xi}_{1}$ $=\left[(3-2\sqrt{3})/6,(3-2\sqrt{3})/6,2\sqrt{3}/3\right],$
$\boldsymbol{\xi}_{2}$ $=\left[(3+2\sqrt{3})/12,(3-\sqrt{3})/6,(3+2\sqrt{3})/12\right],$
$\boldsymbol{\xi}_{3}$ $=\left[(3-\sqrt{3})/6,(3+2\sqrt{3})/12,(3+2\sqrt{3})/12\right],$
$\boldsymbol{\xi}_{4}$ $=\left[2\sqrt{3}/3,(3-2\sqrt{3})/6,(3-2\sqrt{3})/6\right],$
$\boldsymbol{\xi}_{5}$ $=\left[(3+2\sqrt{3})/12,(3+2\sqrt{3})/12,(3-\sqrt{3})/6\right],$
$\boldsymbol{\xi}_{6}$ $=\left[(3-2\sqrt{3})/6,2\sqrt{3}/3,(3-2\sqrt{3})/6\right],$
and radius $r=\|\boldsymbol{v}_{1}-\boldsymbol{v}_{2}\|/(2\sqrt{3}+4)$.
A.3 Derivation of Adaptive Gaussian Covariance
Figure 9: The linear transformation applied on Gaussians for a irregular triangle.
This section presents how to obtain the matrix $\boldsymbol{M}$ used in (10) for a target triangle $(\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3})$.
As shown in Fig. 9, the linear transformation applied on a reference equilateral triangle will result in the target triangle. In the local coordinate system defined in (9), the target triangle is denoted as $\boldsymbol{T}=[\boldsymbol{v}^{\prime}_{1},\boldsymbol{v}^{\prime}_{2},\boldsymbol{v}^{\prime}_{3}]$, and the reference equilateral triangle as $\boldsymbol{E}=[\boldsymbol{v}^{\prime}_{1},\boldsymbol{v}^{\prime}_{2},\hat{\boldsymbol{v}}^{\prime}_{3}]$, where $\hat{\boldsymbol{v}}^{\prime}_{3}=l\cdot[0,\frac{1}{2},\frac{\sqrt{3}}{2}]^{T}$, with $l=\|\boldsymbol{v}_{2}-\boldsymbol{v}_{1}\|$.
By solving the equation $\boldsymbol{M}\boldsymbol{E}=\boldsymbol{T}$, we obtain:
$\boldsymbol{M}=\begin{bmatrix}1&0&0\\ 0&1&\frac{2\boldsymbol{v}^{\prime(2)}_{3}-l}{\sqrt{3}l}\\ 0&0&\frac{2\boldsymbol{v}^{\prime(3)}_{3}}{\sqrt{3}l}\\ \end{bmatrix},$ (13)
where $\boldsymbol{v}^{\prime(i)}_{3}$ denotes the $i$-th component of $\boldsymbol{v}^{\prime}_{3}$.
A.4 More Results
Tab. 7 includes the per-scene results on Mip-NeRF360 dataset [36].